OpenStudy (anonymous):

Very very tired, stuck on test. Answer please? A ball is dropped from a certain height. The function below represents the height f(n), in feet, to which the ball bounces at the nth bounce. f(n) = 9(0.7)^n What does the number 0.7 represent? A. The ball bounces to 30% of its previous height with each bounce. B.The height at which the ball bounces at the nth bounce is 0.3 feet. C.The ball bounces to 70% of its previous height with each bounce. D. The height from which the ball was dropped at the nth bounce is 0.7 feet.

OpenStudy (anonymous):

After no bounces the ball would be 9 * 1 = 9 ft After one bounces the ball would be 9 * .7^1 = 6.3 ft After two bounces the ball would be 9 * .7^2 = 4.41 ft After three bounces the ball would be 9 * .7^3 = 3.087 ft So then what is the .7 doing each bounce?